A cylinder with a 16-cm-diameter piston contains gas at a pressure of 3.0atm.

a) How much force does the gas exert on the piston?

b) How much force does the environment exert on the piston?

c) The gas expands at constant pressure and pushes the piston out 10cm. How much work is done by the environment?

d) How much work is done by the gas?
Answer #1

d=16cm

r=8cm=0.08m

Area A=pi*r^2

=3.14*0.08^2

=2.001*10^-2 m^2

p=3 atm

=3*1.01*10^5 Pascal

a)

Force due to gas is Fg=P*A

=(3*1.01*10^5)*(2.001*10^-2)

=6089.09 N

b)

Force due to environment Fe=P*A

=(3*1.01*10^5)*(2.001*10^-2)

=6089.09 N

C)

l=10 cm=0.1m

now

Work done by the environment on the gas W=P*V

=P*A*l

=6089.09*0.1

=608.909 J (negative, here work is done on the gas)

d)
Work done by the gas w=P*A*l

=6089.09*0.01

W=608.909 J (positive,here work is doing by the gas)